# Simple Self-Organizing Map with Visual Transformer

### v1

@@ IV Conclusion
This work addresses two critical limitations in modern deep learning: Vision Transformers’ (ViTs) lack of inherent inductive biases for limited or small datasets, and Self-Organizing Maps’ (SOMs) suboptimal feature abstraction capabilities. We bridge this gap by proposing ViT-SOM, a framework that synergistically integrates ViTs and SOMs to mutually enhance their strengths. ViT-SOM leverages SOMs to impose topological constraints on ViT embeddings, while ViTs empower SOMs with robust feature extraction in high-dimensional spaces. Empirically validated in both supervised and unsupervised settings, our method demonstrates consistent improvements over baseline and state-of-the-art models, including computationally heavier architectures like ResNet34. Crucially, ViT-SOM achieves this without complex architectural modifications, highlighting the inherent compatibility of these frameworks. By demonstrating the mutual enhancement of ViTs and SOMs, this work encourages broader research in this area.
Figure 4: Decoded prototypes of well-trained ViT-SOM for clustering on MNIST.

### v2

@@ IV Conclusion
This work addresses two critical limitations in modern deep learning: Vision Transformers’ (ViTs) lack of inherent inductive biases for limited or small datasets, and Self-Organizing Maps’ (SOMs) suboptimal feature abstraction capabilities. We bridge this gap by proposing ViT-SOM, a framework that synergistically integrates ViTs and SOMs to mutually enhance their strengths. ViT-SOM leverages SOMs to impose topological constraints on ViT embeddings, while ViTs empower SOMs with robust feature extraction in high-dimensional spaces. Empirically validated in both supervised and unsupervised settings, our method demonstrates consistent improvements over baseline and state-of-the-art models, including computationally heavier architectures like DeiT, Swin, and ResNet34. Crucially, ViT-SOM achieves this without complex architectural modifications, highlighting the inherent compatibility of these frameworks. By demonstrating the mutual enhancement of ViTs and SOMs, this work encourages broader research in this area.